# Two categories of UV-upturn galaxies revealed by semi-analytic models

### v1

@@ Abstract
pubyear: 2024pagerange: Two categories of UV-upturn galaxies revealed by semi-analytic modelsTwo categories of UV-upturn galaxies revealed by semi-analytic models
@@ 1 Introduction
The “UV upturn” or “UV excess” is an unusual phenomenon characterized by excessive far-ultraviolet (FUV) flux observed in some elliptical galaxies and the central bulge of some spiral galaxies, such as M31 (e.g. Rosenfield et al., 2012). This phenomenon was discovered with the first FUV observations by the Orbiting Astronomical Observatory 2 (OAO-2; Code, 1969; Code & Welch, 1979). Early imaging and spectroscopic studies found that the UV excess is extended (e.g. Deharveng et al., 1976), following the profile of visible light (e.g. Bohlin et al., 1985; Welch, 1982), but lacking association with massive stars (e.g. King et al., 1992; Bertola et al., 1995) or star formation activities (e.g. O’Connell et al., 1992). Further spectral and photometric analyses have shown that white dwarfs alone cannot account for the entire UV-upturn phenomenon (Werle et al., 2020). These observations collectively suggest that the UV excess is not due to young massive stars, but rather originates from old stellar populations other than white dwarfs. Additionally, studies examining UV-upturn galaxies across various low-redshift environments have found no dependence of UV excess on cluster environment, group multiplicity, or satellite velocity dispersion. This suggests that the UV upturn is driven by processes internal to the galaxies themselves (e.g. Yi et al., 2011; Loubser & Sánchez-Blázquez, 2011; Boissier et al., 2018; Ali et al., 2018c, 2019; Phillipps et al., 2020).
@@
Theoretically, extreme horizontal branch (EHB) stars and their hot post-HB descendants, including post-early asymptotic giant branch (AGB) stars and AGB-manqué stars, have been proposed as key candidates responsible for the UV excess in old, long-lived stellar populations (Greggio & Renzini, 1990, 1999). EHB stars are core-helium-burning stars characterized by extremely thin hydrogen envelopes ($M_{\rm env}<0.02\,{\rm M}_{\odot}$), with temperatures $T_{\rm e}\sim 25,000\,{\rm K}$ and masses of approximately $0.5\,{\rm M}_{\odot}$. They fall below the main sequence of massive stars on the Hertzsprung-Russell (H-R) diagram and are particularly energetic in the FUV band. Given that $10-20\%$ of evolving stars pass through the EHB channels, the observed FUV luminosities can be reasonably explained by EHB stars (Greggio & Renzini, 1999; Brown et al., 1995). The detection of EHB and post-EHB stars in M32 by the Hubble Space Telescope provided direct evidence for the EHB origin of the UV excess (Brown et al., 2000b, 2008); for a comprehensive review, see O’Connell (1999).
@@
2 Semi-analytic Model
@@
As mentioned, UV excess is observed in ellipticals or galactic bulges. In order to facilitate a more reasonable comparison between model and observation, we only consider the “bulge region” of each model galaxy. The “bulge region” of a galaxy is defined as the area within twice the half-mass radius of the stellar bulge, which is formed mainly through galaxy mergers in gabe. Assuming the density profile of the galactic disk to be exponential, we calculate the volume fraction of this bulge region over the whole disk and assign the corresponding fraction of disk component to the bulge. In this way, we will compare different properties of the bulge region of model galaxies with observations, and the model galaxies mentioned hereafter are represented by their bulge regions. Note that a model galaxy defined in this way is naturally a composite of a bulge with or without a disk. The ones without disks are analogous to fully quenched elliptical galaxies in observations, while those with disks can be compared with elliptical galaxies with certain levels of star formation activities or with the bulges of spiral galaxies in observations. We only consider modeled galaxies whose bulge region is more massive than $10^{9.5}\,{\rm M}_{\odot}$.
@@ 3.1 Colors of simple stellar populations
We first examine the magnitudes and color indices in UV and optical bands for Simple Stellar Populations (SSPs) across different Evolutionary Population Synthesis (EPS) models. The left panels in Figure 1 illustrate the evolution of magnitude in the Far UV (FUV), Near UV (NUV), and $r$ band for a $1\,{\rm M}_{\odot}$ SSP with solar metallicity ($Z=0.02$), for all six EPS models. The models show negligible differences in both NUV and $r$ magnitude when stellar age exceeds $\sim 0.01$ Gyr. In contrast, significant variations among the models are observed in the FUV magnitude for SSPs older than $\sim 0.3$ Gyr, where YII-bin and YII-ehb models are more luminous than YII-sin and galaxev in the FUV band, attributable to the increasing impact of binary interactions. These interactions, such as mass transfer and merger, give rise to hotter stellar objects, predominantly hot subdwarfs and blue stragglers [^5] Blue stragglers are stars lying on the main sequence in the color-magnitude diagram of globular clusters while beyond the turning-point (Sandage, 1953). They are believed to be produced from collisions between low-mass stars and mass transfer in close binaries (e.g., Pols & Marinus, 1994; Chen & Han, 2004; Hurley et al., 2005)., enhancing the FUV luminosity around $1\,{\rm Gyr}$ compared to single star scenarios. At ages larger than $\sim 3$ Gyr, Extreme Horizontal Branch (EHB) stars produced through the three formation channels of Han et al. (2002, 2003) play a more significant role, lowering the FUV magnitude by an additional magnitude, potentially accounting for the UV-upturn phenomenon observed in elliptical galaxies, as discussed in Han et al. (2007). The bpass models share a similar evolution with that of yunnan-ii, but are about $1-2$ magnitudes fainter in the FUV band at large ages. Due to the lack of formation channels of EHB stars of Han et al. (2002, 2003), the single and binary models of bpass gradually converge with each other at $\sim 3$ Gyr.
@@
Figure 4: Left: Redshift evolution of rest-frame NUV-$r$ colors of modeled quenched galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$) run with various EPS models as shown in the legend. Thick solid lines and dotted lines are the median values, and shaded regions indicate the $25\%$ and $75\%$ quartiles. Dashed lines mark the range corresponding to an additional 1.5 times the interquartile range. The YII-ehb and BPASS-sin models are fully illustrated as thick solid lines with corresponding distributions as the lower and upper limits of model predictions, respectively; and the median values of other EPS models are marked as dotted lines for clarity. Box plots are the observational results of ETGs in 12 clusters in Ali et al. (2021), whose boxes and whiskers share the same definition as our shaded regions and dashed lines, respectively. Circles are observed outliers beyond the whisker range in each cluster, and diamonds are stacks in two clusters at high redshifts. Right: The same as the left panel, but for the observed NUV-i colors. Data points are the observational results of field red galaxies of De Propris et al. (2022) and Atlee et al. (2009), and their error bars illustrate the root mean squares of the mean values from 100 bootstrapping realizations within each magnitude/redshift bin.
@@ 3.2 Colors and redshift evolution of model galaxies
We now extend the analysis to higher redshifts. The left panel of Figure 4 shows the redshift evolution of the rest-frame NUV-$r$ color for quenched galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$) in different EPS models. The YII-ehb and BPASS-sin models yield the bluest and reddest NUV-$r$ colors, respectively; therefore, these two models are fully illustrated as thick solid lines with corresponding distributions to indicate the plausible range predicted by different EPS models. The median values of other EPS models are marked as dotted lines for clarity. For comparison, symbols with error bars show ETGs within 12 galaxy clusters at $1.00.3<z<$ 1.0 from the Hubble Legacy Archive (Ali et al., 2021). While this study does not aim to quantitatively compare our models with observations, it is noteworthy that the model derived with YII-ehb agrees quite well with the observational results, in terms of both the median and the scatter of NUV-$r$ at a given redshift. However, it is important to note that the differences across different EPS models are comparable to or even smaller than the inherent scatter of galaxies.
In the right panel of Figure 4, we compare the predicted redshift evolution of the "observed" NUV-$i$ color of quenched model galaxies with field red galaxies from De Propris et al. (2022). Following De Propris et al. (2022), we use the rest-frame FUV-$r$ color from the EPS models as analogues of the observed NUV-$i$ colors [^7] We find the measured NUV-$i$ colors are in fact $\sim 1-2$ bluer than the rest-frame FUV-$r$ colors due to the broader bandwidth of the NUV band (Morrissey et al., 2005). We have corrected this effect for each modeled galaxy according to its redshift and mean mass-weighted age and metallicity, by utilizing spectra of each EPS model.. As observed, the YII-ehb model exhibits marginal agreement with the upper limits of the observational results, whereas other EPS models yield excessively redder colors. The variations of the rest-frame FUV-$r$ color among the EPS models are significantly larger than those of the rest-frame NUV-$r$ color seen in the left panel, especially at high redshift. This is consistent with what we have seen from Figure 1, where the FUV-involved colors are more powerful than NUV-involved colors in distinguishing different EPS models at old ages. Ali et al. (2021) and De Propris et al. (2022) suggested that the He-rich single star model provided the best explanation for their observations. Our analysis further suggests that the binary star model YII-ehb is also compatible with current observations in terms of the redshift evolution of quenched galaxies at $z\lesssim 1$.
@@ 3.3 Influence of stellar initial mass function
$\Phi(\log_{\rm 10}{(m_{\rm 1})})\propto\begin{cases}\left(m_{\rm 1}/0.6\right)^{-\Gamma_{\rm bi}}&\text{(for $m_{\rm 1}>0.6\,{\rm M}_{\odot}$)}\\ p\left(m_{\rm 1}\right)&\text{(for $0.2\leq m_{\rm 1}\leq 0.6\,{\rm M}_{\odot}$),}\\ 1&\text{(for $m_{\rm 1}<0.2\,{\rm M}_{\odot}$)}\end{cases}$ (1)
@@ 3.4 Influence of dust attenuation
There are four parameters in this parameterization: the slope $\delta$ accounts for additional wavelength dependence that diverges from the Calzetti00 curve; $E_{\rm b}$, $\lambda_{\rm 0}$ and $\Delta\lambda$ denote the amplitude, central wavelength, and full width at half maximum (FWHM) of the UV bump, respectively (Fitzpatrick & Massa, 1990). The latter two parameters are set to average values, specifically $\lambda_{\rm 0}=2175$ and $\Delta\lambda=350$ (Noll et al., 2009a; Seaton, 1979; Kriek & Conroy, 2013), and we investigate the parameter space of the slope $\delta$ and the UV bump amplitude $E_{\rm b}$. We consider two sets of models of dust attenuation curves with varying parameters $\delta$ and $E_{\rm b}$ as listed in  LABEL:tab:dust_models. The first set is aimed to isolate the effects of UV bump strength, with a fixed slope of $\delta=0$ but three different values of $E_{\rm b}$. The second set of models are assumed to follow the observed relation of $E_{\rm b}=-1.9\delta+0.85$ from Kriek & Conroy (2013).

### v2

@@ Abstract
pubyear: 2025pagerange: Two categories of UV-upturn galaxies revealed by semi-analytic modelsA.2
@@ 1 Introduction
The “UV upturn” or “UV excess” is an unusual phenomenon characterized by excessive far-ultraviolet (FUV) flux observed in some elliptical galaxies and the central bulge of some spiral galaxies, such as M31 (e.g. Rosenfield et al., 2012). This phenomenon was discovered with the first FUV observations by the Orbiting Astronomical Observatory 2 (OAO-2; Code, 1969; Code & Welch, 1979). Early imaging and spectroscopic studies found that the UV excess is spatially extended (e.g. Deharveng et al., 1976, 1980, 1982), following the profile of visible light (e.g. Welch, 1982; Bohlin et al., 1985), but lacking association with massive stars (e.g. King et al., 1992; Bertola et al., 1995) or star formation activities (e.g. O’Connell et al., 1992). Further spectral and photometric analyses have shown that white dwarfs alone cannot account for the entire UV-upturn phenomenon (Werle et al., 2020). These observations collectively suggest that the UV excess is not due to young massive stars, but rather originates from old stellar populations other than white dwarfs. Additionally, studies examining UV-upturn galaxies across various low-redshift environments have found no dependence of UV excess on cluster environment, group multiplicity, or satellite velocity dispersion. This suggests that the UV upturn is driven by processes internal to the galaxies themselves (e.g. Yi et al., 2011; Loubser & Sánchez-Blázquez, 2011; Boissier et al., 2018; Ali et al., 2018c, 2019; Phillipps et al., 2020).
@@
Theoretically, extreme horizontal branch (EHB) stars and their hot post-HB descendants, including post-early asymptotic giant branch (AGB) stars and AGB-manqué stars, have been proposed as key candidates responsible for the UV excess in old, long-lived stellar populations (Greggio & Renzini, 1990, 1999). EHB stars are core-helium-burning stars characterized by extremely thin hydrogen envelopes ($M_{\rm env}<0.02\,{\rm M}_{\odot}$), with temperatures $T_{\rm e}\sim 25,000\,{\rm K}$ and masses of approximately $0.5\,{\rm M}_{\odot}$. They fall below the main sequence of massive stars on the Hertzsprung-Russell (H-R) diagram and are particularly energetic in the FUV band. Given that $10-20\%$ of evolving stars pass through the EHB channels, the observed FUV luminosities can be reasonably explained (Greggio & Renzini, 1999; Brown et al., 1995). The detection of EHB and post-EHB stars in M32 by the Hubble Space Telescope provided direct evidence for the EHB origin of the UV excess (Brown et al., 2000b, 2008); for a comprehensive review, see O’Connell (1999). We note that, in this work, we do not meticulously distinguish the effect of EHB stars with that of their descendants. Hereafter the term “EHB stars” represents the sum of them for simplicity.
@@
2 Method
2.1 Semi-analytic Model
@@
As mentioned, UV excess is observed in ellipticals or galactic bulges. In order to facilitate a more reasonable comparison between model and observation, we only consider the “bulge region” of each model galaxy. The “bulge region” of a galaxy is defined as the area within twice the half-mass radius of the stellar bulge, which is formed mainly through galaxy mergers in gabe. Assuming the density profile of the galactic disk to be exponential, we calculate the volume fraction of this bulge region over the whole disk and assign the corresponding fraction of disk component to the bulge. In this way, we will compare different properties of the bulge region of model galaxies with observations, and the model galaxies mentioned hereafter are represented by their bulge regions. Note that a model galaxy defined in this way is naturally a composite of a bulge with or without a disk. The ones without disks are analogous to fully quenched elliptical galaxies in observations, while those with disks can be compared with elliptical galaxies with certain levels of star formation activities or with the bulges of spiral galaxies in observations. We only consider modeled galaxies whose bulge region is more massive than $10^{9.5}\,{\rm M}_{\odot}$, to reflect the fact that UV-upturn galaxies are found mainly in massive galaxies (e.g. Le Cras et al., 2016; Akhil et al., 2024; Martocchia et al., 2025) as well as to match the stellar mass lower limit we apply on the observational sample in subsection 3.2.
@@
2.2 Definition of “UV-upturn”
In this work, we adopt the UV-upturn classification scheme of Yi et al. (2011), in which a UV-upturn galaxy is defined to have NUV-$r>5.4$, FUV-NUV $<0.9$, and FUV-$r<6.6$. These photometric criteria were designed to select quenched galaxies with a reliable upturn feature:
  • NUV-$r>5.4$ is adopted to seperate fully-quenched galaxies from those exhibiting relatively prominent residual star formation (RSF) activities. As shown by previous studies of GALEX observations, the NUV-$r$ color presents a bimodal distribution and is a good indicator of RSF activities in galaxies (e.g. Kaviraj et al., 2007; Salim et al., 2007; Salim, 2014). This criterion was originally proposed by Yi et al. (2005), and the exact value adopted varies from $5.0$ to $5.5$ (e.g. 5.0, Yi et al. 2005, Jeong et al. 2009, Salim 2014, Crossett et al. 2017; 5.2, Phillipps et al. 2020; 5.4, Schawinski et al. 2007, Yi et al. 2011; 5.5, Kaviraj et al. 2007). Here we use a relatively strict value of $5.4$, to better focus on quenched galaxies. It is noteworthy that even though galaxies with prominent RSF activities can be removed by this single color cut, the ones with arbitrarily small amounts of RSF can never be totally ruled out, as discussed in Phillipps et al. (2020).
  • FUV-NUV $<0.9$ is a criterion to recognize the upturn feature towards shorter wavelengths, based on the fact that a flat UV spectrum $F(\lambda)$ typically has a FUV-NUV color of 0.9 (Yi et al., 2011). However, as suggested in Smith et al. (2012) and Phillipps et al. (2020), the UV properties of quenched galaxies actually behave in a “continuous” way and their FUV-NUV colors span a relatively wide range (e.g. 0.7-2.1 in Phillipps et al., 2020). Besides, UV SEDs of different quenched ETGs or different regions within ETGs are found to be better fitted by hot components with different temperatures (e.g. $10,000-21,000\,{\rm K}$ in Ali et al. 2018a; $25,000-40,000\,{\rm K}$ in Martocchia et al. 2025), rather than a single temperature; and even some of the FUV-NUV$>0.9$ galaxies prefer a model with an extra hot component, e.g. M86 in Martocchia et al. (2025). These results suggest that the mechanism which induces the upturn phenomenon can not be easily picked out by a single color cut. Such a “continuous” behavior also presents in this work (check the ranges of FUV-NUV colors of old stellar populations in Figure 1 and Figure 2). Nevertheless, we adopt the FUV-NUV $<0.9$ criterion in this work to focus on the most extreme part of this continuum – quenched galaxies which indeed have an upturn feature. Note that the value of 0.9 for FUV-NUV may slightly vary depending on the EPS model assumed. As this threshold is used to select a subpopulation within a continuous distribution, its tiny variance will have no influence on our conclusions.
  • FUV-$r$ $<6.6$ reflects the relatively high flux ratio between FUV to optical bands, as observed in UV-upturn galaxies. For galaxies without RSF activities, the FUV flux is a better representative of the hot evolved component than the NUV flux, which is additionally contributed by metal line blanketing from the main sequence turn-off (MSTO) stars (e.g. Donas et al., 2007; Smith et al., 2012; Martocchia et al., 2025). Therefore, FUV-$r$ color works here as a pure comparison between hot evolved stars and normal cool old stars. Furthermore, by demanding a relatively strong FUV emission, the reliability of the selected upturn feature can be ensured. The value of 6.6 was calculated for a typical quenched galaxy with FUV-NUV$=0.9$ in Yi et al. (2011), and it is also model dependent. Our test analysis shows that this model dependence has no significant influence on our conclusions, though.
Therefore, the UV-upturn classification scheme of Yi et al. (2011) as described above is an empirical, model-dependent, and relatively strict definition. We adopt this definition to focus on the most typical UV-upturn galaxies, where the mechanism behind the UV-upturn phenomenon plays a more influential role. In fact, the definition or measurement of the UV-upturn phenomenon varies from one paper to another. Different definitions may result in different samples and lead to different results. In subsection 4.1, we will come back to this point and discuss on the influence of the definition of UV-upturn on our results.
@@ 3.1 Colors of simple stellar populations
We first examine the magnitudes and color indices in UV and optical bands for Simple Stellar Populations (SSPs) across different Evolutionary Population Synthesis (EPS) models. The left panels in Figure 1 illustrate the evolution of absolute magnitude in the Far UV (FUV), Near UV (NUV), and $r$ band for a $1\,{\rm M}_{\odot}$ SSP with solar metallicity ($Z=0.02$), for all six EPS models. The models show negligible differences in both NUV and $r$ magnitude when stellar age exceeds $\sim 0.01$ Gyr. In contrast, significant variations among the models are observed in the FUV magnitude for SSPs older than $\sim 0.3$ Gyr, where YII-bin and YII-ehb models are more luminous than YII-sin and galaxev in the FUV band, attributable to the increasing impact of binary interactions. These interactions, such as mass transfer and merger, give rise to hotter stellar objects, predominantly hot subdwarfs and blue stragglers [^5] Blue stragglers are stars lying on the main sequence in the color-magnitude diagram of globular clusters while beyond the turning-point (Sandage, 1953). They are believed to be produced from collisions between low-mass stars and mass transfer in close binaries (e.g., Pols & Marinus, 1994; Chen & Han, 2004; Hurley et al., 2005)., enhancing the FUV luminosity around $1\,{\rm Gyr}$ compared to single star scenarios. At ages larger than $\sim 3$ Gyr, Extreme Horizontal Branch (EHB) stars produced through the three formation channels of Han et al. (2002, 2003) play a more significant role, lowering the FUV magnitude by an additional magnitude, potentially accounting for the UV-upturn phenomenon observed in elliptical galaxies, as discussed in Han et al. (2007). The bpass models share a similar evolution with that of yunnan-ii, but are about $1-2$ magnitudes fainter in the FUV band at large ages. Due to the lack of formation channels of EHB stars of Han et al. (2002, 2003), the single and binary models of bpass gradually converge with each other at $\sim 3$ Gyr.
@@
Figure 4: Left: Redshift evolution of rest-frame NUV-$r$ colors of modeled quenched galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$) run with various EPS models as shown in the legend. Thick solid lines and dotted lines are the median values, and shaded regions indicate the $25\%$ and $75\%$ quartiles. Dashed lines mark the range corresponding to an additional 1.5 times the interquartile range. The YII-ehb and BPASS-sin models are fully illustrated as thick solid lines with corresponding distributions as the lower and upper limits of model predictions, respectively; and the median values of other EPS models are marked as dotted lines for clarity. Box plots are the observational results of ETGs in 12 clusters in Ali et al. (2021), whose boxes and whiskers share the same definition as our shaded regions and dashed lines, respectively. Circles are observed outliers beyond the whisker range in each cluster, and diamonds are stacks in two clusters at high redshifts. The lower panel shows the differences between various EPS models and YII-ehb. Right: Redshift evolution of the rest-frame FUV-$r$ colors (for various EPS models) and the observed NUV-$i$ colors (for observational data points). The lines are the same as in the left panel. Data points are the observational results of field red galaxies of De Propris et al. (2022) and Atlee et al. (2009), and their error bars illustrate the root mean squares of the mean values from 100 bootstrapping realizations within each magnitude/redshift bin.
@@ 3.2 Colors and redshift evolution of model galaxies
We now extend the analysis to higher redshifts. The left panel of Figure 4 shows the redshift evolution of the rest-frame NUV-$r$ color for quenched galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$) in different EPS models. The YII-ehb and BPASS-sin models yield the bluest and reddest NUV-$r$ colors, respectively; therefore, these two models are fully illustrated as thick solid lines with corresponding distributions (shaded regions: $25\%-75\%$ quartiles; dashed lines: an additional 1.5 times the interquartile range) to indicate the plausible range predicted by different EPS models. The median values of other EPS models are marked as dotted lines for clarity. For comparison, box plots show the distribution of ETGs within 12 galaxy clusters at $1.00.3<z<$ 1.0 from the Hubble Legacy Archive (Ali et al., 2021), whose boxes and whiskers share the same definition as our shaded regions and dashed lines, respectively. Circles are observed outliers beyond the whisker range in each cluster, and diamonds are stacks in two clusters at high redshifts. While this study does not aim to quantitatively compare our models with observations, it is noteworthy that the model derived with YII-ehb agrees quite well with the observational results, in terms of both the median and the scatter of NUV-$r$ at a given redshift. However, it is important to note that the differences across different EPS models are comparable to or even smaller than the inherent scatter of galaxies.
In the right panel of Figure 4, we compare the predicted redshift evolution of the "observed" NUV-$i$ color of quenched model galaxies with field red galaxies from De Propris et al. (2022). Following De Propris et al. (2022), we use the rest-frame FUV-$r$ color from the EPS models as analogues of the observed NUV-$i$ colors given their redshift ranges [^7] We find the measured NUV-$i$ colors are in fact $\sim 1-2$ magnitudes bluer than the rest-frame FUV-$r$ colors due to the broader bandwidth of the NUV band (Morrissey et al., 2005). We have corrected this effect for each modeled galaxy according to its redshift and mean mass-weighted age and metallicity, by utilizing spectra of each EPS model.. As observed, the YII-ehb model exhibits marginal agreement with the upper limits of the observational results, whereas other EPS models yield excessively redder colors. The variations of the rest-frame FUV-$r$ color among the EPS models are significantly larger than those of the rest-frame NUV-$r$ color seen in the left panel, especially at high redshift. This is consistent with what we have seen from Figure 1, where the FUV-involved colors are more powerful than NUV-involved colors in distinguishing different EPS models at old ages. Ali et al. (2021) and De Propris et al. (2022) suggested that the He-rich single star model provided the best explanation for their observations. Our analysis further suggests that the binary star model YII-ehb is also compatible with current observations in terms of the redshift evolution of quenched galaxies at $z\lesssim 1$.
@@ 3.3 Influence of stellar initial mass function
$\Phi(\log_{\rm 10}{(m_{\rm 1})})\propto\begin{cases}(m_{\rm 1}/0.6)^{-\Gamma_{\rm bi}}&\text{(for $m_{\rm 1}>0.6\,{\rm M}_{\odot}$)}\\ p(m_{\rm 1})&\text{(for $0.2\leq m_{\rm 1}\leq 0.6\,{\rm M}_{\odot}$),}\\ 1&\text{(for $m_{\rm 1}<0.2\,{\rm M}_{\odot}$)}\end{cases}$ (1)
@@
The redshift evolution of the rest-frame NUV-$r$ and FUV-$r$ colors of modeled quenched galaxies are presented in Appendix A.2, for EPS models with different IMF slopes. Consistent with the result obtained above, variation in IMF slope has a negligible effect on the UV-to-optical colors in old stellar populations.
@@ 3.4 Influence of dust attenuation
There are four parameters in this parameterization: the slope $\delta$ accounts for additional wavelength dependence that diverges from the Calzetti00 curve; $E_{\rm b}$, $\lambda_{\rm 0}$ and $\Delta\lambda$ denote the amplitude, central wavelength, and full width at half maximum (FWHM) of the UV bump, respectively (Fitzpatrick & Massa, 1990). The latter two parameters are set to average values, specifically $\lambda_{\rm 0}=2175$ and $\Delta\lambda=350$ (Noll et al., 2009a; Seaton, 1979; Kriek & Conroy, 2013), and we investigate the parameter space of the slope $\delta$ and the UV bump amplitude $E_{\rm b}$. We consider two sets of models of dust attenuation curves with varying parameters $\delta$ and $E_{\rm b}$ as listed in LABEL:tab:dust_models. The first set is aimed to isolate the effects of UV bump strength, with a fixed slope of $\delta=0$ but three different values of $E_{\rm b}$. The second set of models are assumed to follow the observed relation of $E_{\rm b}=-1.9\delta+0.85$ from Kriek & Conroy (2013). Graphs of these curves are presented in Appendix A.1.
The redshift evolution of the UV-to-optical colors of quenched galaxies with various dust attenuation models are discussed in Appendix A.1. We find that different dust models merely have a negligible influence, with $\Delta$(NUV-$r$) $\lesssim 0.10$ and $\Delta$(FUV-$r$) $\lesssim 0.15$, as these quenched galaxies are naturally devoid of dust.
3.5 Joint dependence on $F_{\rm b}$, $F_{\rm y}$, $Z_{\rm*}$ and dust